OpenStudy (anonymous):

which equation is an identity? a.8-(5x+2)=-5x-6 b.7z+10-z=8z-2(z-5) c.8m-4=5m+8-m d.6y+5=6y-5 which equation has no solution? a.7v+2=8v-3 b.3x-5=3x+8-3 c.4y+5=4y-6 d.7z+6=-7z-5

OpenStudy (anonymous):

3+6z=13+6z a. z=-5/6 b.z=2 2/3 c.many solutions d.no solution -6p+7=3(2p-3)-4(-10+4p) a. p=6 b.p=5 c.p=7 d.p=12

OpenStudy (_julianna_):

I know that the first question that you posted up first is D :)

OpenStudy (anonymous):

thanks